OpenStudy (anonymous):

Find dy/dx using the chain rule. y = (ax-b)^4

OpenStudy (anonymous):

do you know the chain rule?

OpenStudy (xapproachesinfinity):

\(\large \frac{d}{dx}(f^n)=nf'f^{n-1}\)

OpenStudy (aum):

\[ y = (ax-b)^4 \\ \frac{dy}{dx} = \frac{d}{dx}(ax-b)^4 = 4(ax-b)^3 * \frac{d}{dx}(ax-b) = ? \]